OpenStudy (anonymous):

(3x+y)/y=6/5. Solve for x and y.

OpenStudy (australopithecus):

\[\frac{3x+y}{y} = \frac{6}{5}\] \[y*\frac{3x+y}{y} = \frac{6}{5}*y\] \[3x+y= \frac{6y}{5}\] \[3x= \frac{6y}{5} - y\] \[3x= (\frac{6}{5} - 1)y\] \[\frac{3x}{(\frac{6}{5} - 1)}= \frac{(\frac{6}{5} - 1)y}{(\frac{6}{5} - 1)}\] \[\frac{3x}{(\frac{6}{5} - 1)}= y\]

OpenStudy (anonymous):

i was looking for numerical values

OpenStudy (australopithecus):

This is a function I cannot give you a numerical value if you do not give me a value for x or y

OpenStudy (australopithecus):

I did solve for y though, based on what I applied to solve for y you should be able to solve for x